# Morpheus: Benchmarking Physical Reasoning of Video Generative Models with Real Physical Experiments

### v1

@@ Abstract
Recent advances in image and video generation raise hopes that these models possess world modeling capabilities—the ability to generate realistic, physically plausible videos. This could revolutionize applications in robotics, autonomous driving, and scientific simulation. However, before treating these models as world models, we must ask: Do they adhere to physical conservation laws? To answer this, we introduce Morpheus, a benchmark for evaluating video generation models on physical reasoning. It features 80 real-world videos capturing physical phenomena, guided by conservation laws. Since artificial generations lack ground truth, we assess physical plausibility using physics-informed metrics evaluated with respect to infallible conservation laws known per physical setting, leveraging advances in physics-informed neural networks and vision-language foundation models. Our findings reveal that even with advanced prompting and video conditioning, current models struggle to encode physical principles despite generating aesthetically pleasing videos. All data, leaderboard, and code are open-sourced at: https://physics-from-video.github.io/morpheus-bench/
Benchmark Real-world ground-through Quantitative evaluation Initial condition grounding Physical laws evaluation
VideoCon-Physics [1]  (only text)
VAMP [2]  (only text)
PhyGenBench [3]  (only text)
Kang et al. [4]  (1 or 3 frames)
COSMOS [5]  (image and video)
Physics-IQ [6]  (image and video)
Morpheus (ours)  (image and video)
Table 1: Comparison of physics-based video understanding benchmarks. Symbols: = supported, = not supported
@@ 1 Introduction
Video generative models (VGMs) such as SORA [7], Veo2 [8], and COSMOS [5] have taken the world by storm, building upon remarkable advances in image generative models [9, 10, 11, 12], and achieving unprecedented levels of visual fidelity and realism.
These developments have not only pushed the boundaries of visual aesthetics but have also inspired the community to envision video generative models as potential world models [13, 5]. A world model, in this context, is more than just a system for generating frames, however; it is a model capable of understanding and predicting the dynamics, causal interactions, and underlying mechanisms of the physical world. Accurately benchmarking the physical dynamics of video generation is a critical requirement —and the focus of this work— toward adopting them potentially as world models.
Evaluating physical dynamics of generated video is far from straightforward. When describing physical systems,
Figure 1: Physical Evaluation of VGMs for the holonomic pendulum experiment using Dynamical and Physical Invariance scores computed from the extracted trajectories.
[…]
@@
We propose Morpheus, a novel benchmarking framework designed to evaluate the physical reasoning capabilities of video generative models using real-world physical experiments. The key idea behind Morpheus is to map video recordings of physical events—whether generated by models or recorded from real experiments—into a common physical representation that can be analyzed and compared. Leveraging advances in zero-shot object segmentation, object tracking, and physics-informed neural networks (PINNs) [14, 15], our framework a) fits to the video dynamics the ODE dynamics that should be governing the underlying system, and b) extracts standardized physical measurements, such as velocity and acceleration from video data, which should conform to conservation laws. By collecting measurements from both real physical videos and generated ones, and comparing their summary statistics with respect to governing ODEs and physical invariants, Morpheus enables fair and systematic benchmarking of physical invariants, such as the conservation of energy or momentum, without requiring explicit ground truth data.
We make four contributions toward benchmarking the physical reasoning capabilities of video generative models.
First, we introduce Morpheus, the first benchmark, including a public leaderboard, using real-world physical experiments to systematically evaluate physical reasoning based explicitly on physical invariants (section 3).
Second, we propose a novel framework that combines physics-informed deep learning with advanced computer vision techniques to enable coarse- and fine-grained analysis of physical phenomena (section 4).
Third, we evaluate state-of-the-art video generative models on Morpheus generating over $9000$ videos, including CogVideoX [16], PyramidalFlow [17], LTX-Video [18] and COSMOS [5], and show that while these models excel in visual aesthetics, they fall short in modeling real-world physical dynamics (section 5).
Finally, we highlight key limitations and provide actionable insights for improving the physical reasoning capabilities of video generation models.
@@ 2 Related work
Benchmarking video generation models have evolved to include comprehensive evaluation frameworks that assess multiple dimensions of video quality, temporal coherence, and alignment with prompts. Approaches like EvalCrafter [19], VBench [20], VBench++ [21], AIGCBench [22], and TC-Bench [23] emphasize diverse metrics to evaluate visual fidelity, motion smoothness, spatial consistency, and temporal dynamics. For example, EvalCrafter [19] uses metrics like Motion-Aware Consistency (MAC) and Scene Change Consistency (SCC) to assess the smoothness and natural progression of motion, while VBench introduces metrics for spatial relationships and subject identity consistency to evaluate logical scene composition. Despite the breadth of these benchmarks, they primarily concentrate on perceptual and semantic aspects of video generation, whereas Morpheus focuses on physical plausibility of the generated videos.
@@
Recent advances in evaluating physical plausibility in video generation have employed both human assessments [1] and automated approaches leveraging vision-language models (VLMs) [1, 3] and object tracking metrics [2, 6, 5]. Notable frameworks include VideoCon-Physics [1], PhyGenBench [3] and PhysBench [24], which utilize VLMs to assess adherence to physical law prompts; VAMP [2], which quantifies motion characteristics through acceleration and velocity variance; and Physics-IQ [6] and COSMOS [5], which compare object masks between generated and real-world videos. Kang et al. [4] used the PHYRE simulator [25] to fine-tune VGM on synthetic 2D data, facilitating out-of-distribution and combinatorial generalization evaluation.
Despite addressing diverse physical phenomena, these benchmarks have significant limitations. VLM and human evaluations often identify physical deviations categorically, like noting gravity violations without quantifying them. Moreover, VLM can hallucinate [26] and miss subtle physical inconsistencies. On the other side, object tracking metrics are often based on simulated data [5, 25] and assume that modeled processes should be deterministic and predictable [4, 6]. These limitations highlight the critical need for more robust, interpretable benchmarks that can quantitatively evaluate physical realism by precisely measuring how well-generated videos preserve physical invariants and adhere to specific physical laws.
@@
There is progress for learning conservation laws from trajectories [27], and equation discovery in hybrid dynamic systems  [28]. Mechanistic Neural Networks (MechNN) [29] are able to learn governing Ordinary Differential Equations (ODEs) from data, while Mechanistic PDE Networks [30] can learn Partial Differential Equations (PDEs). On the other hand, to compare the theoretical prediction with input data, PINNs [14, 15], who integrate physical equations in the loss function, can help identify possible physical factors causing errors (such as unmodeled friction, air drag, etc.) because it is able to learn corrections to make the predictions closer to the actual observed values.
@@ 3 Morpheus Benchmark
To rigorously examine the discrepancies in adherence to physical laws within generated videos, we propose the Morpheus benchmark.
3.1 Methodology for creating the dataset
@@
We created a dataset of real-world recordings of specific physical phenomena, focusing on fundamental aspects of Newtonian mechanics. Recordings were conducted under controlled laboratory conditions, allowing us to systematically vary initial parameters and capture repeatable scenarios. By operating in this rigorously controlled setting, we can isolate and test adherence to specific physical laws – such as the periodic dynamics of a harmonic pendulum – rather than merely assessing overall visual plausibility. This sets our dataset apart from previous works that often focus on uncontrolled, general-purpose video data [6, 4], allowing for a more precise and targeted evaluation of physical consistency.
Physical experiments.
We recorded a set of six core experiments, each highlighting specific physical principles:
  1. 1.
    Falling ball: A ball dropped from rest until it makes impact with the surface, used to test uniform gravitational acceleration and energy conservation.
  2. 2.
    Bouncing ball: A ball observed from the moment it first impacts the surface until it rebounds and impacts again, testing gravitational acceleration and energy conservation in a more challenging setting.
  3. 3.
  4. 4.
  5. 5.
    Non-holonomic pendulum: A ball suspended by a string, showcasing approximate energy conservation and approximate harmonic motion.
  6. 6.
Figure 2: Different types of discarded generated videos: (left) A video showing the disappearance of the orange ball during fall; (middle) A video illustrating generation of multiple non-holonomic pendulums; (right) A video in which the double pendulum does not move.
Physical initial conditions.
For each experiment, we varied specific initial conditions: the initial speed for the falling ball, the angle and initial speed for the projectile motion, and the angle for the pendulums. This diverse collection ensures robust coverage of dynamic behaviors, enabling thorough evaluation of generated videos against real-world physical phenomena. To maintain consistency in these initial parameters, we employed an actuator – namely, a robotic controller – ensuring accurate and repeatable setups. We recorded 5–7 videos per initial condition setting, resulting in approximately 10-20 videos per experiment. Complete dataset statistics and the photos of the experiments can be found in App. A.
We use the recorded dataset images in two ways mainly in our experiments. First, the initial frames of the real-world videos serve as prompts to set the initial conditions for the conditional video generation models. This ensures that the generated sequences begin from the same starting point as the real-world experiments.
@@
Second, we use real-world videos as a “gold standard” to validate that our evaluation metrics work as intended. By analyzing the metrics on these ground-truth recordings, we demonstrate the reliability of our approach and establish an upper bound on performance, representing the precision with which we can measure adherence to physical laws. In essence, the metrics computed on real-world videos provide a baseline for how closely any generative model can align with physical principles.
@@
To structurally analyze the videos, we extract the trajectories of the objects by applying promptable visual segmentation. These trajectories comprise 2D coordinates of the recognized objects through time and are used for further analysis with our physical metrics. We describe the trajectory extraction in Sec. 3.3.
Figure 3: The video generation (upper) and the trajectory extraction pipeline (lower). We use the first frame (or multiple frames in case of video conditioning) of the real-world recording of an experiment, as well as the textual description, as a prompt for a Video Generation Model (VGM). Next, we extract object trajectories for both real-world and generated videos using trajectory extraction pipeline, consisting of Segmentation and Tracking (SAM2), and relative depth estimation (DepthAnything V2) and postprocessing.
@@ 3.2 Prompting methods
Figure 4: Evaluation of trajectories extracted from real and VGMs videos using Physical Invariance score derived from physical invariants.
The physical dynamics of a scene are fundamentally determined by its initial conditions, which include the positions, velocities, and geometric constraints (e.g. shapes, rigid body connections) of all objects at the outset. In the context of generative models, these initial conditions are set through prompting. There are three main approaches to prompting: a) textual prompts, b) single-image prompts, and c) video (or multi-frame) prompts. Each provides a different level of control over the generation process.
Textual prompts offer minimal control. They can suggest the general behavior or nature of a scene - such as a ball rolling or falling - but lack the precision to define exact starting positions, velocities, or trajectories. Single-image prompts improve upon this by allowing the initial positions of objects to be visually specified. However, they still fall short in terms of velocity and motion details. Only video prompts, which incorporate sequences of frames, grant the ability to set both initial positions and velocities of all currently visible objects. This introduces a gradation in the level of control: from the broad suggestions of textual prompts to the detailed specifications of video prompts.
With this gradation in mind, we investigate how different levels of control affect the physical realism of the generated samples. We explore both textual prompt enhancement and various multi-frame prompting for models capable of leveraging these features (e.g. [5, 16]), allowing us to examine the relationship between input precision and output physical fidelity. Following the approach of [16], simple scene descriptions can be enhanced with the advanced capabilities of a VLM [31] in creating rich and descriptive prompts along when provided with a certain instruction template in either a zero-shot or few-shot fashion. As not all VGMs provide a prompt upsampler along the model, we rely on the ChatGLM family of models [31], while for COSMOS [5] we use their own devised [32, 33] upsamplers. In our evaluation, we still try to create a highly descriptive prompt with an emphasis on capturing the physical motion, and the upsampler allows us to bring the distribution of the textual prompt during inference closer to the one used during training.

### v2

@@ Abstract
Recent advances in image and video generation raise hopes that these models possess world modeling capabilities—the ability to generate realistic, physically plausible videos. This could revolutionize applications in robotics, autonomous driving, and scientific simulation. However, before treating these models as world models, we must ask: Do they adhere to physical laws? Current evaluation methods rely on subjective judgments or trajectory matching, limiting their usage for physical reasoning estimation, where many generations could be physically plausible. Thus, we introduce Morpheus, a new benchmark for evaluating video generation models on physical reasoning. It features 130 real-world videos capturing physical phenomena, guided by conservation laws. Using those as conditioning for video generation, we assess physical plausibility using physics-informed metrics evaluated with respect to infallible conservation laws known per physical setting, leveraging advances in physics-informed neural networks and vision-language foundation models. Our findings reveal that even with advanced prompting and video conditioning, current models struggle to encode physical principles despite generating aesthetically pleasing videos. All data, leaderboard, and code are open-sourced at: https://physics-from-video.github.io/morpheus-bench/
@@ 1 Introduction
Video generative models (VGMs) such as SORA (Brooks et al., 2024a), COSMOS (Agarwal et al., 2025), and Veo3 (Veo-Team et al., 2024) have taken the world by storm, building upon remarkable advances in image generative models (Ramesh et al., 2021; Saharia et al., 2022; Podell et al., 2023; Yu et al., 2023), and achieving unprecedented levels of visual fidelity and realism. These developments have not only pushed the boundaries of visual aesthetics but have also inspired the community to envision video generative models as potential world models (Cho et al., 2024; Agarwal et al., 2025). A world model, in this context, is more than just a system for generating frames, however; it is a model capable of understanding and predicting the dynamics, causal interactions, and underlying mechanisms of the physical world. Accurately benchmarking the physical dynamics of video generation is a critical requirement —and the focus of this work— toward adopting them potentially as world models.00footnotetext: Signs $\ast$, $\dagger$, $\ddagger$ denote equal first, second, and last author contributions correspondingly. Perhaps the most daring challenge from an AI perspective is the need for physical dynamics evaluation that goes beyond visual verification. Current methods use either a) human or VLM judgement (e.g., “does this video of an object falling look legitimate?”)  (Bansal et al., 2024; Meng et al., 2024), or b) visual or geometric plausibility (e.g., “is the generated scene visually consistent through time?”)  (Agarwal et al., 2025) or c) comparsion with one possible future using trajectory matching (e.g., comparing the generated locations of a projectile with one possible ground-truth locations) (Kang et al., 2024; Motamed et al., 2025)). While human-based evaluations (potentially distilled in VLMs) are useful for initial screening of generated videos, they do not provide sufficient quantitative and objective evidence of physical plausibility. Trajectory matching evaluation can yield false negatives in cases where the generated video is physically plausible, but the object’s trajectory deviates from a particular ground-truth trajectory due to unobserved initial conditions or other visually hidden factors, such as an object’s mass or friction, that affect the object’s motion.
A deeper physical understanding requires assessing whether a generated video preserves physical invariants and physical dynamics that govern the underlying system. For instance, in many systems, quantities such as total energy must remain constant as the system evolves, providing opportunities for quantitative benchmarks of physical plausibility. In addition, we can test whether an object’s trajectory is consistent with the set of trajectories permitted by the governing physical laws by evaluating the fit of the observed dynamics to the governing equations of motion. Using these quantitative evaluations, we can design systematic benchmarks that reveal whether video generative models (VGMs) truly capture the dynamics of the physical world or simply produce visually plausible approximations.
Figure 1: Comparison of evaluation methods for video generative models. a) Human or VLM-based judgments provide only qualitative and subjective assessments of physical plausibility. b) Trajectory matching compares generated and ground-truth paths but may misclassify physically valid trajectories. For example, for projectile motion, many parabolic trajectories are physically plausible when VGMs are conditioned only on an image, as object velocity cannot be estimated from it. c) Our proposed framework, Morpheus, evaluates generated videos via physics-informed scores, testing both conservation of physical invariants and consistency with governing equations of motion.
@@
We propose Morpheus, a novel physics-informed benchmarking framework designed to evaluate the physical reasoning capabilities of video generative models using real-world physical experiments. The key idea behind Morpheus is to map video recordings of physical events—whether generated by models or recorded from real experiments—into a common physical representation that can be analyzed and compared. Leveraging advances in zero-shot object segmentation, object tracking, and physics-informed neural networks (PINNs) (Cuomo et al., 2022; He et al., 2023), our framework a) fits to the video dynamics the ODE that governs the underlying system, and b) extracts standardized physical measurements, such as velocity and acceleration from video data, which should conform to conservation laws. By collecting measurements from both real physical videos and generated ones, and comparing their summary statistics with respect to governing ODEs and physical invariants, Morpheus enables fair and systematic benchmarking of physical invariants, such as the conservation of energy or momentum, without requiring or relying on explicit ground truth data.
This work makes three contributions toward rigorous evaluation of physical reasoning in video generative models. First, we introduce Morpheus, the first benchmark to systematically evaluate physical reasoning based explicitly on physical invariants (Section 3) using real-world physical experiments to ground VGM’s generations in a controllable setting. Second, we propose a novel framework that combines physics-informed deep learning with advanced computer vision techniques to enable coarse- and fine-grained analysis of physical phenomena (Section 4). Third, we evaluate state-of-the-art video generative models on Morpheus generating over $9000$ videos, including CogVideoX (Yang et al., 2024b), PyramidalFlow (Jin et al., 2024), LTX-Video (HaCohen et al., 2024), Wan2.1 (Wan et al., 2025b), COSMOS (Agarwal et al., 2025) and Veo3 (Veo-Team et al., 2024), and show that while the best of these models excel in visual aesthetics, they fall short in modeling real-world physical dynamics (Section 5).
@@ 2 Related work
Benchmarking video generation models have evolved to include comprehensive evaluation frameworks that assess multiple dimensions of video quality, temporal coherence, and alignment with prompts. Approaches like EvalCrafter (Liu et al., 2024a), VBench (Huang et al., 2024a), VBench++ (Huang et al., 2024b), AIGCBench (Fan et al., 2024), and TC-Bench (Feng et al., 2024) emphasize diverse metrics to evaluate visual fidelity, motion smoothness, spatial consistency, and temporal dynamics. For example, EvalCrafter (Liu et al., 2024a) uses metrics like Motion-Aware Consistency (MAC) and Scene Change Consistency (SCC) to assess the smoothness and natural progression of motion, while VBench introduces metrics for spatial relationships and subject identity consistency to evaluate logical scene composition. Despite the breadth of these benchmarks, they primarily concentrate on perceptual and semantic aspects of video generation, whereas Morpheus focuses on physical plausibility of the generated videos.
Figure 2: The overview of Morpheus benchmark. Video augmentation and generation (upper) and the trajectory extraction pipelines (lower). We start with augmenting recorded videos with realistic style transfer, based on object masks. Next, we use the first frame (or multiple frames in case of video conditioning) of the obtained videos, as well as the textual description, as a prompt for a VGM. After this, we extract object trajectories for both real-world and generated videos using the trajectory extraction pipeline, including trajectory tracking and discarding unreliable trajectories. Finally, we evaluate Morpheus scores for all videos with valid trajectories.
@@
Recent advances in evaluating physical plausibility in video generation have employed both human assessments (Bansal et al., 2024) and automated approaches leveraging vision-language models (VLMs) (Bansal et al., 2024; Meng et al., 2024) as well as object tracking metrics (Wang et al., 2024b; Motamed et al., 2025; Agarwal et al., 2025) (see Table 2 for comparison). Notable frameworks include VideoCon-Physics (Bansal et al., 2024), PhyGenBench (Meng et al., 2024) and PhysBench (Chow et al., 2025), which utilize VLMs to assess adherence to physical law prompts; VAMP (Wang et al., 2024b), which quantifies motion characteristics through acceleration and velocity variance; and Physics-IQ (Motamed et al., 2025) and COSMOS (Agarwal et al., 2025), which compare object masks between generated and real-world videos. Kang et al. (Kang et al., 2024) used the PHYRE simulator (Bakhtin et al., 2019) to fine-tune VGM on synthetic 2D data, facilitating out-of-distribution and combinatorial generalization evaluation.
Despite addressing diverse physical phenomena, these benchmarks have significant limitations. VLM and human evaluations often identify physical deviations categorically, like noting gravity violations without quantifying them. Moreover, VLM can hallucinate (Li et al., 2023) and miss subtle physical inconsistencies (Chow et al., 2025). On the other side, object tracking metrics are often based on simulated data (Agarwal et al., 2025; Bakhtin et al., 2019) and assume that modeled processes should be deterministic and predictable (Kang et al., 2024; Motamed et al., 2025). These limitations highlight the critical need for more robust, interpretable benchmarks that can quantitatively evaluate physical realism by precisely measuring how well-generated videos preserve physical invariants and adhere to specific physical laws.
@@
There is progress for learning conservation laws from trajectories (Liu & Tegmark, 2021), and equation discovery in hybrid dynamic systems  (Liu et al., 2024b). Mechanistic Neural Networks (MechNN) (Pervez et al., 2024) are able to learn governing ODEs from data, while Mechanistic PDE Networks (Pervez et al., 2025) can learn Partial Differential Equations (PDEs). On the other hand, to compare the theoretical prediction with input data, PINNs (Cuomo et al., 2022; He et al., 2023), which integrate physical equations in the loss function, help identify possible physical factors causing errors (such as unmodeled friction, air drag, etc.) because it is able to learn corrections to make the predictions closer to the actual observed values.
Figure 3: Examples of physical experiments included in the Morpheus benchmark, illustrating both different dynamics and variations in object types. Top row (left to right): falling ball, projectile motion, holonomic pendulum, sliding, falling apple, and double pendulum. Bottom row (left to right): collision, falling tape, rolling can, spring, rolling orange, and bouncing.
@@ 3 Morpheus Benchmark
To rigorously examine discrepancies in adherence to physical laws within generated videos, we propose the Morpheus benchmark, consisting of a dataset for controlled conditioning, evaluation scores, and an analysis of state-of-the-art models’ performance.
Dataset methodology
@@
We created a dataset of real-world videos of specific physical phenomena, focusing on capturing fundamental aspects of Newtonian mechanics. Videos were recorded under controlled laboratory conditions, allowing us to systematically vary initial parameters and capture repeatable scenarios. By operating in this rigorously controlled setting, we can isolate and test adherence to specific physical laws – such as the periodic dynamics of a harmonic pendulum – rather than merely assessing overall visual plausibility. This sets our dataset apart from previous works that often focus on uncontrolled, general-purpose video data (Motamed et al., 2025; Kang et al., 2024), allowing for a more precise and targeted evaluation of physical consistency.
We recorded a set of nine core physical experiments with 10-20 videos in each experiment, each highlighting different physical principles such as gravity (falling, projectile motion, and bouncing), periodic movements (spring and holonomic pendulum), friction and normal forces (incline sliding and rolling), and more complex dynamical systems such as multi-body collisions and double pendulum (See Fig. 3 and App. A for dataset an illustration and description). For each experiment, we varied the initial conditions (recording 5–7 videos per condition), such as speed for falling, angle and speed for projectile motion, and angle for pendulums. This diverse collection ensures robust coverage of dynamic behaviors, enabling thorough evaluation of generated videos against real-world physical phenomena. The initial video frame(s) are used as conditioning to guide VGM’s sampling/generation process, ensuring the generated sequences start from the same conditions as the real experiments.
Visually diverse conditioning for robust VGMs evaluation.
To obtain more robust evaluations, it is important to study the performance of VGMs under diverse natural initializations of the same physical process. This reduces sensitivity to the specific setup of the originally recorded experiments. Thus, we augment the initial videos with visually diverse yet reliably generated variants. In particular, we use the video-to-video transfer method (Alhaija et al., 2025) with object masks extracted from recorded videos to obtain diverse and visually realistic videos following an additional text prompt for adaptation. While changing the semantics and appearance of the objects, the generated videos strictly follow the provided object masks and are reliable augmentations for conditioning in image-to-video and video-to-video generation. To ensure their quality, we filter augmented initializations through manual inspection for visual and physical plausibility. Thus, the original initializations are expanded 10-fold by generating 3 variations with 3 stylization prompts. For more details, we refer the reader to App. B.
Metrics validation.
@@
We use real-world videos as a “gold standard” to validate that our evaluation metrics work as intended. By analyzing the metrics on these ground-truth recordings, we demonstrate the reliability of our approach and establish an upper bound on performance, representing the precision with which we measure adherence to physical laws. In essence, the metrics computed on real-world videos provide a baseline for how closely any generative model can align with physical principles.
@@
To structurally analyze the videos, we extract the trajectories of the objects by applying promptable video segmentation. These trajectories comprise 2D coordinates of the recognized objects through time and are used for further analysis with our physical metrics (see Sec. 3.2 for details).
@@ 3.1 Prompting methods
The physical dynamics of a scene are determined by its initial conditions: object positions, velocities, and geometric constraints (e.g., shapes, rigid connections). In generative models, these conditions are set through prompting, which can take three forms: a) textual prompts, b) single-image prompts, and c) video (multi-frame) prompts, providing a different level of control over generation. Textual prompts offer only broad control, suggesting behaviors (e.g., rolling, falling) without precise states. Single-image prompts improve precision by fixing initial locations, but lack motion detail. Only video prompts specify both positions and velocities, providing the highest control.
With this gradation in mind, we investigate how different levels of control affect the physical realism. We explore both textual prompt enhancement and various multi-frame prompting for models capable of leveraging these features (e.g.  (Agarwal et al., 2025; Yang et al., 2024b)), allowing us to examine the relationship between input precision and output physical fidelity. Following (Yang et al., 2024b), we use a VLM (Zeng et al., 2024) to expand simple scene descriptions into richer prompts via instruction templates in zero- or few-shot settings. As not all VGMs provide their own prompt upsampler, we rely on the ChatGLM family of models (Zeng et al., 2024), while for COSMOS-variants (Agarwal et al., 2025) and WAN-2.1 (Wan et al., 2025a) we use their own devised (NVIDIA, 2024; Agrawal et al., 2024; Wang et al., 2024a) upsamplers respectively. In our evaluation, we create descriptive prompts with an emphasis on physical motion, and the upsampler brings the inference-time prompt distribution closer to that used during training.